the manager of a branch office of banco mexicali observed that during peak hours an average of 25 customers arrives per hour and that there is an average of 10 customers in the branch office at any time. How long does the average customer spend wainting in the line and being serviced.

The average time a customer spends in the branch office including service is.... hours.( enter your response rounded to two decimal places)
Answer #1

Let the bank operates between 10 AM to 5 PM

Lunch break = 30 min from 1PM to 1:30 PM

Total working hrs.= 7-.5= 6.5 hrs

The peak period of a bank is generally in the first half= 1PM-10AM = 3 hrs

No. of customers in peak hr= 3*25= 75

Time for serving =3 hrs

At say 11 AM no. of customers in bank= 10

While between 10-11 AM no of customers entered = 25-10= 15

Total time taken for 15 customers = 60 min

Time taken for each customer = 60/15= 4 min